Suppose y is directly proportional to x
(a) If y=8 when x= 13, find k so that y=kx.
(b) Then use y=kx to find y when x=3

Answer by RAY100(1637)

y=kx
.
8=k*13,,,,,given (13,8)
.
8/13 =k
.
y=8/13 * x = (8/13) * 3 = 24 /13, when x =3